# The norms for symmetric and antisymmetric tensor products of the weighted shift operators

### v1

@@ 2 Preliminaries
$\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}=\left\{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}:\exists\,\pi\in\Sigma_{n}\text{ and }\exists\,1\leq m\leq n~{}s.t.~{}\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}.$
@@
$\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}=\left\{\mathfrak{j}\in\mathbb{N}^{n}:|\mathfrak{j}|=k+S_{\mathfrak{l}},\exists\,\pi\in\Sigma_{n}~{}s.t.~{}\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}$
@@
$\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}=\left\{\mathfrak{j}\in\mathbb{N}^{n}:\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}},\exists\,\pi\in\Sigma_{n}\text{ and }\exists\,1\leq m\leq n~{}s.t.~{}\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}.$
@@
#$^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq\ ^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}$ (2.35)
@@
$=$ $\bigcup_{m=1}^{n}\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}.$
Suppose that ${}^{\#}\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}\neq\emptyset.$ Then for $\mathfrak{j}\in\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}$, there exists a $\pi\in\Sigma_{n},$ such that $l_{\pi(m)}\neq 0,$ and $\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon$, i.e.
@@
Recall that $h=\max\left\{{\left|l_{1}\right|,\cdots,\left|l_{n}\right|}\right\}$ is introduced in (2.17). Hence by (2.37) and (2.38), for $\mathfrak{j}\in\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}$, $j_{m}\leq N+h+1.$ It follows that for every $m\in\left\{1,2,\cdots,n\right\}$,
$\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}\subseteq\bigcup\limits_{j_{m}=0}^{N+h+1}\left\{\mathfrak{j}\in\mathbb{N}^{n}:|\mathfrak{j}|=k+S_{\mathfrak{l}}\right\}.$ (2.39)
It is not hard to verify that
{${}^{\#}\left\{\mathfrak{j}\in\mathbb{N}^{n}:|\mathfrak{j}|=k+S_{\mathfrak{l}}\right\}\leq\left(k+S_{\mathfrak{l}}\right)^{n-2},$ (2.40)
@@
${}^{\#}\check{B}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon,m}\leq(N+h+1)\left(k+S_{\mathfrak{l}}\right)^{n-2}.$ (2.41)
@@
${}^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq n(N+h+1)\left(k+S_{\mathfrak{l}}\right)^{n-2}\leq Ck^{n-2}.$
@@
$\bigcup_{j=0}^{\mathscr{M}}E_{k+S_{\mathfrak{l}},j}=\left\{\left\langle\mathfrak{j}\right\rangle\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}$
@@
{${}^{\#}\left\{\left\langle\mathfrak{j}\right\rangle\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}=P(k+S_{\mathfrak{l}},n).$
@@
#$^{\#}E_{k+S_{\mathfrak{l}},\mathscr{M}}=P(k+S_{\mathfrak{l}},n)-\sum\limits_{j=0}^{\mathscr{M}-1}\ {}^{\#}E_{k+S_{\mathfrak{l}},j}.$ (2.42)
@@
${}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ =#$=\ ^{\#}E_{k+S_{\mathfrak{l}},\mathscr{M}}-\ ^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ (2.43)
@@
For every $10\leq j\leq\mathscr{M}-$ 1 and sufficiently large $k$, by (2.22), (2.12) and Lemma 2.2, there exists a constant $M>0$ such that
@@
Hence for every $10\leq j\leq\mathscr{M}-$ 1,
@@
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k+S_{\mathfrak{l}},n)}=\lim\limits_{k\to\infty}\frac{P(k+S_{\mathfrak{l}},n)-\sum\limits_{j=0}^{\mathscr{M}-1}\ {}^{\#}E_{k+S_{\mathfrak{l}},j}-\ ^{\#}\check{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}{P(k+S_{\mathfrak{l}},n)}=1.$
@@
For $k\geq 0$ and $0\leq i_{1}\leq\cdots\leq i_{n}$ satisfying $|\mathfrak{i}|=k,$ we have that for $\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}(0<\varepsilon<1)$ and $\pi\in N_{\mathfrak{j},\mathfrak{i}},$ it holds
@@
$\left\langle\mathfrak{i}+\mathfrak{l}_{\pi}\right\rangle=\left\langle\mathfrak{j}\right\rangle.$
Then for every fixed $1\leq a\leq n$, there exists a $1\leq b\leq n$, such that $i_{a}+l_{\pi(a)}=j_{b}.$ Choose a $\tau_{0}\in\Sigma_{n}$ such that $\tau_{0}(b)=\pi(a)$. Hence from $\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}$ and (2.33), we have
@@
$\left(1-\varepsilon\right)^{|l_{\pi(a)}|}<\left|\beta_{i_{a},l_{\pi(a)}}\right|=\begin{cases}\left|\alpha_{i_{a}-1}\cdots\alpha_{i_{a}+l_{\pi(a)}}\right|,&-l_{\pi(a)}>0,\\ 1,&-l_{\pi(a)}=0,\\ \left|\alpha_{i_{a}}\cdots\alpha_{i_{a}+l_{\pi(a)}-1}\right|,&-l_{\pi(a)}<0.\end{cases}$ (2.48)
@@
$P_{\mathfrak{i},k,t}=\left\{\pi\in\Sigma_{n}:\left\langle\mathfrak{i}+\mathfrak{l}_{\pi}\right\rangle\in E_{k+S_{\mathfrak{l}},t}\right\},\quad 0\leq t\leq\mathscr{M}.$ (2.50)
@@
$P_{\mathfrak{i},k,t}=\bigsqcup\limits_{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},t}}N_{\mathfrak{j},\mathfrak{i}}.$ (2.51)
@@
$\Sigma_{n,\mathfrak{i}}^{\prime}=\mathop{\bigsqcup}_{t=1}^{\mathscr{M}}\bigsqcup\limits_{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},t}}N_{\mathfrak{j},\mathfrak{i}}.$ (2.52)
@@ 3.1 Proof of “$(1)\Leftrightarrow(2)$
$\left\|{1\over\mu}S_{\alpha,l_{1}}\odot\cdots\odot{1\over\mu}S_{\alpha,l_{n}}\right\|={1\over|\mu|^{|\mathfrak{l}|}}\left\|S_{\alpha,{l_{1}}}\odot S_{\alpha,{l_{2}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right\|.$
@@
Without loss of generality assume $l>0$. Since $1=\lim\limits_{i\to\infty}\left|\alpha_{i}\right|\geq\left|\alpha_{m}\right|$ for any $m\in\mathbb{N}$, it follows that for any $01>\varepsilon>$ 0, there exists an $N>0$ such that for all $i>N$,
@@
$\left(S_{\alpha}^{l}\odot S_{\alpha}^{l}\odot\cdots\odot S_{\alpha}^{l}\right)\left(e_{\langle\mathfrak{i}\rangle}\right)=\left(\prod\limits_{j=1}^{n}\prod\limits_{k=1}^{l}\alpha_{i_{j}+k-1}\right)\left(e_{i_{1}+l}\odot e_{i_{2}+l}\odot\cdots\odot e_{i_{n}+l}\right).$
@@
For $0\leq i_{1}\leq\cdots\leq i_{n}$ with $|\mathfrak{i}|=k$, let $a_{\langle\mathfrak{i}\rangle}=\frac{1}{\sqrt{P(k,n)}}$. Obviously $\sum\limits\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{\langle\mathfrak{i}\rangle}e_{\langle\mathfrak{i}\rangle}$ is a unit vector. To calculate the norm of $S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}$, we will estimate the norms of $S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{\langle\mathfrak{i}\rangle}e_{\langle\mathfrak{i}\rangle}\right)$. By the definition of $\Sigma_{n,\mathfrak{i}}^{\prime}$ in (2.49), it is easy to see that for $\pi\in\Sigma_{n}\backslash\Sigma_{n,\mathfrak{i}}^{\prime},$
@@
$\left(S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right)\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{\langle\mathfrak{i}\rangle}\frac{e_{\langle\mathfrak{i}\rangle}}{\left\|e_{\langle\mathfrak{i}\rangle}\right\|}$ (3.5)
$=$ $\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\frac{a_{\langle\mathfrak{i}\rangle}}{n!}\sum\limits_{\pi\in\Sigma_{n,\mathfrak{i}}^{\prime}}\frac{\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}e_{\langle\mathfrak{i}+\mathfrak{l}_{\pi}\rangle}}{\left\|e_{\langle\mathfrak{i}\rangle}\right\|}$
$=$ $\sum\limits_{\langle\mathfrak{j}\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$
$~{}+\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{\langle\mathfrak{j}\rangle\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$
$=$ $\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$
$~{}+\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\check{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$
$~{}+\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}.$
For $\mathfrak{j}\in\mathbb{N}^{n}$ satisfying $\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$, by (2.26), we have
[…]
@@
$\left\|e_{\langle\mathfrak{j}\rangle}\right\|=\frac{1}{\sqrt{n!}}.$ (3.6)
@@
$\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\tilde{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle},$
$\sum\limits_{\left\langle\mathfrak{j}\right\rangle\in\check{A}_{k+\left|\mathfrak{l}\right|,\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{\langle\mathfrak{i}\rangle}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{\langle\mathfrak{i}\rangle}\right\|}\right)e_{\langle\mathfrak{j}\rangle}$

### v2

@@ 2 Preliminaries
$\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}=\left\{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}:\exists\,\pi\in\Sigma_{n}\text{ and }\exists\,1\leq m\leq n~s.t.~\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}.$
@@
$\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}=\left\{\mathfrak{j}\in\mathbb{N}^{n}:|\mathfrak{j}|=k+S_{\mathfrak{l}},\exists\,\pi\in\Sigma_{n}~s.t.~\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}$
@@
$\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}=\left\{\mathfrak{j}\in\mathbb{N}^{n}:[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}},\exists\,\pi\in\Sigma_{n}\text{ and }\exists\,1\leq m\leq n~s.t.~\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon\right\}.$
@@
$^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq\ ^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}^{\prime}$ (2.35)
@@
$=$ $\bigcup_{m=1}^{n}\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}.$
Suppose that $^{\#}\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}\neq\emptyset$. Then for $\mathfrak{j}\in\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}$, there exists a $\pi\in\Sigma_{n},$ such that $l_{\pi(m)}\neq 0,$ and $\Gamma_{j_{m},-l_{\pi(m)}}\leq 1-\varepsilon$, i.e.
@@
Recall that $h=\max\left\{{\left|l_{1}\right|,\cdots,\left|l_{n}\right|}\right\}$ is introduced in (2.17). Hence by (2.37) and (2.38), for $\mathfrak{j}\in\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}$, $j_{m}\leq N+h+1.$ It follows that for every $m\in\left\{1,2,\cdots,n\right\}$,
$\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}\subseteq\bigcup\limits_{j_{m}=0}^{N+h+1}\left\{\mathfrak{j}\in\mathbb{N}^{n}:|\mathfrak{j}|=k+S_{\mathfrak{l}}\right\}.$ (2.39)
It is not hard to verify that for every fixed $0\leq t\leq N+h+1$,
$^{\#}\left\{\mathfrak{j}\in\mathbb{N}^{n}:j_{m}=t,|\mathfrak{j}|=k+S_{\mathfrak{l}}\right\}\leq\left(k+S_{\mathfrak{l}}+1\right)^{n-2},$ (2.40)
@@
$^{\#}\check{B}_{k+S_{\mathfrak{l}},\varepsilon,m}\leq(N+h+2)\left(k+S_{\mathfrak{l}}+1\right)^{n-2}.$ (2.41)
@@
$^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}\leq n(N+h+2)\left(k+S_{\mathfrak{l}}+1\right)^{n-2}\leq Ck^{n-2}.\qed$
@@
$\bigcup_{j=0}^{\mathscr{M}}E_{k+S_{\mathfrak{l}},j}=\left\{[\mathfrak{j}]\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}$
@@
$^{\#}\left\{[\mathfrak{j}]\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{j}\right|=k+S_{\mathfrak{l}}\right\}=P(k+S_{\mathfrak{l}},n).$
@@
$^{\#}E_{k+S_{\mathfrak{l}},\mathscr{M}}=P(k+S_{\mathfrak{l}},n)-\sum\limits_{j=0}^{\mathscr{M}-1}\ {}^{\#}E_{k+S_{\mathfrak{l}},j}$. (2.42)
@@
${}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ $=\ ^{\#}E_{k+S_{\mathfrak{l}},\mathscr{M}}-\ ^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ (2.43)
@@
For every $0\leq j\leq\mathscr{M}-1$ and sufficiently large $k$, by (2.22), (2.12) and Lemma 2.2, there exists a constant $M>0$ such that
@@
Hence for every $0\leq j\leq\mathscr{M}-1$,
@@
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k+S_{\mathfrak{l}},n)}=\lim\limits_{k\to\infty}\frac{P(k+S_{\mathfrak{l}},n)-\sum\limits_{j=0}^{\mathscr{M}-1}\ {}^{\#}E_{k+S_{\mathfrak{l}},j}-\ ^{\#}\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k+S_{\mathfrak{l}},n)}=1.$
@@
$\lim\limits_{k\to\infty}\frac{{}^{\#}\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}{P(k,n)}=1.\qed$
@@
For $k\geq 0$ and $0\leq i_{1}\leq\cdots\leq i_{n}$ satisfying $|\mathfrak{i}|=k,$ we have that for $\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}(0<\varepsilon<1)$ and $\pi\in N_{\mathfrak{j},\mathfrak{i}},$ it holds
@@
$\mathfrak{i}+\mathfrak{l}_{\pi}]=[\mathfrak{j}].$
Then for every fixed $1\leq a\leq n$, there exists a $1\leq b\leq n$, such that $i_{a}+l_{\pi(a)}=j_{b}.$ Choose a $\tau_{0}\in\Sigma_{n}$ such that $\tau_{0}(b)=\pi(a)$. Hence from $\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}$ and (2.33), we have
@@
$\left(1-\varepsilon\right)^{|l_{\pi(a)}|}<\left|\beta_{i_{a},l_{\pi(a)}}\right|=\begin{cases}\left|\alpha_{i_{a}-1}\cdots\alpha_{i_{a}+l_{\pi(a)}}\right|,&-l_{\pi(a)}>0,\\ 1,&-l_{\pi(a)}=0,\\ \left|\alpha_{i_{a}}\cdots\alpha_{i_{a}+l_{\pi(a)}-1}\right|,&-l_{\pi(a)}<0.\end{cases}$
The lemma is proved. ∎
@@
$P_{\mathfrak{i},k,t}=\left\{\pi\in\Sigma_{n}:[\mathfrak{i}+\mathfrak{l}_{\pi}]\in E_{k+S_{\mathfrak{l}},t}\right\},\quad 0\leq t\leq\mathscr{M}.$ (2.49)
@@
$P_{\mathfrak{i},k,t}=\bigsqcup\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},t}}N_{\mathfrak{j},\mathfrak{i}}.$ (2.50)
@@
$\Sigma_{n,\mathfrak{i}}^{\prime}=\mathop{\bigsqcup}_{t=1}^{\mathscr{M}}\bigsqcup\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},t}}N_{\mathfrak{j},\mathfrak{i}}.$ (2.51)
@@ 3.1 Proof of $(1)\Leftrightarrow(2)$
$\left\|S_{\frac{\alpha}{\mu},l_{1}}\odot\cdots\odot S_{\frac{\alpha}{\mu},l_{n}}\right\|={1\over|\mu|^{|\mathfrak{l}|}}\left\|S_{\alpha,{l_{1}}}\odot S_{\alpha,{l_{2}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right\|.$
@@
Without loss of generality assume $l>0$. Since $1=\lim\limits_{i\to\infty}\left|\alpha_{i}\right|\geq\left|\alpha_{m}\right|$ for any $m\in\mathbb{N}$, it follows that for any $1>\varepsilon>0$, there exists an $N>0$ such that for all $i>N$,
@@
$\left(S_{\alpha}^{l}\odot S_{\alpha}^{l}\odot\cdots\odot S_{\alpha}^{l}\right)\left(e_{[\mathfrak{i}]}\right)=\left(\prod\limits_{j=1}^{n}\prod\limits_{k=1}^{l}\alpha_{i_{j}+k-1}\right)\left(e_{i_{1}+l}\odot e_{i_{2}+l}\odot\cdots\odot e_{i_{n}+l}\right).$
@@
For $0\leq i_{1}\leq\cdots\leq i_{n}$ with $|\mathfrak{i}|=k$, let $a_{[\mathfrak{i}]}=\frac{1}{\sqrt{P(k,n)}}$. Obviously $\sum\limits\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{[\mathfrak{i}]}\frac{e_{[\mathfrak{i}]}}{\left\|e_{[\mathfrak{i}]}\right\|}$ is a unit vector. To calculate the norm of $S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}$, we will estimate the norms of $S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{[\mathfrak{i}]}\frac{e_{[\mathfrak{i}]}}{\left\|e_{[\mathfrak{i}]}\right\|}\right)$. By the definition of $\Sigma_{n,\mathfrak{i}}^{\prime}$ in (2.48), it is easy to see that for $\pi\in\Sigma_{n}\backslash\Sigma_{n,\mathfrak{i}}^{\prime},$
@@
$\left(S_{\alpha,{l_{1}}}\odot\cdots\odot S_{\alpha,{l_{n}}}\right)\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}a_{[\mathfrak{i}]}\frac{e_{[\mathfrak{i}]}}{\left\|e_{[\mathfrak{i}]}\right\|}$ (3.5)
$=$ $\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\frac{a_{[\mathfrak{i}]}}{n!}\sum\limits_{\pi\in\Sigma_{n,\mathfrak{i}}^{\prime}}\frac{\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}e_{[\mathfrak{i}+\mathfrak{l}_{\pi}]}}{\left\|e_{[\mathfrak{i}]}\right\|}$
$=$ $\sum\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$
$~+\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$
$=$ $\sum\limits_{[\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$
$~+\sum\limits_{[\mathfrak{j}]\in\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$
$~+\sum\limits_{t=1}^{\mathscr{M}-1}\sum\limits_{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},t}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}.$
For $\mathfrak{j}\in\mathbb{N}^{n}$ satisfying $\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$, by (2.26), we have
$j_{k_{1}}\neq j_{k_{2}},\ 1\leq k_{1}\neq k_{2}\leq n,$
@@
$\left\|e_{[\mathfrak{j}]}\right\|=\frac{1}{\sqrt{n!}}.$ (3.6)
@@
$\sum\limits_{[\mathfrak{j}]\in\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]},$
$\sum\limits_{[\mathfrak{j}]\in\check{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}}\left(\sum\limits_{0\leq i_{1}\leq\cdots\leq i_{n},|\mathfrak{i}|=k}\sum\limits_{\pi\in N_{\mathfrak{j},\mathfrak{i}}}\frac{a_{[\mathfrak{i}]}\beta_{i_{1},l_{\pi(1)}}\cdots\beta_{i_{n},l_{\pi(n)}}}{n!\left\|e_{[\mathfrak{i}]}\right\|}\right)e_{[\mathfrak{j}]}$